# Recursive subhomogeneity of orbit-breaking subalgebras of $\mathrm{C}^{*}$-algebras associated to minimal homeomorphisms twisted by line bundles

### v1

@@ 8.0.1. RSH structure of $\mathcal{O}(\mathcal{E}_{Y})$
    Let $(X,\alpha)$ be a minimal dynamical system. In [3], it is shown that crossed product $C(X)\rtimes_{\alpha}\mathbb{Z}$ has stable rank one, with no restrictions an $(X,\alpha)$. In particular, $C(X)\rtimes_{\alpha}\mathbb{Z}$ may have stable rank one even if it is not $\mathcal{Z}$-stable. This is shown by exhibiting a so-called diagonal subhomogeneous (DSH) algebra decomposition for orbit-breaking subalgebras when $Y\subset X$ is closed with non-empty interior. A DSH algebra is a special type of RSH algebra where the gluing maps between stages of the RSH decomposition are assumed to be diagonal (see [3, Definition 2.3] for a precise definition). As above, if $y$ is a single point, then this implies $\mathrm{C}^{*}(X,C_{0}(X\setminus\{y\})u)\subset C(X)\rtimes_{\alpha}\mathbb{Z}$ is approximately DSH. They show that a simple inductive limit of DSH algebras with diagonal connecting maps has stable rank one [3, Theorem 3.30]. Thus $\mathrm{C}^{*}(X,C_{0}(X\setminus\{y\})u)$ has stable rank one, and as it is a centrally large subalgebra, this passes to $C(X)\rtimes_{\alpha}\mathbb{Z}$. The RSH constructions in this paper will not be DSH algebra in general. This follows from [3, Corollary 2.22]. Nevertheless, our RSH constructions should be amenable to similar techniques. The question of stable rank of $\mathrm{C}^{*}$-algebras associated to homeomorphisms twisted by (non-trivial) line bundles is the subject of future investigations.

### v2

@@ 8.0.1. An RSH structure for $\mathcal{O}(\mathcal{E}_{Y})$
    Let $(X,\alpha)$ be a minimal system. In [3], it is shown that crossed product $C(X)\rtimes_{\alpha}\mathbb{Z}$ has stable rank one, with no restrictions an on $(X,\alpha)$. In particular, $C(X)\rtimes_{\alpha}\mathbb{Z}$ may have stable rank one even if it is not $\mathcal{Z}$-stable. This is shown by exhibiting a so-called diagonal subhomogeneous (DSH) algebra decomposition for orbit-breaking subalgebras when $Y\subset X$ is closed with non-empty interior. A DSH algebra is a special type of RSH algebra where the gluing maps between stages of the RSH decomposition are assumed to be diagonal (see [3, Definition 2.3] for a precise definition). As above, if $y$ is a single point, then this implies $\mathrm{C}^{*}(X,C_{0}(X\setminus\{y\})u)\subset C(X)\rtimes_{\alpha}\mathbb{Z}$ is approximately DSH. They show that a simple inductive limit of DSH algebras with diagonal connecting maps has stable rank one [3, Theorem 3.30]. Thus $\mathrm{C}^{*}(X,C_{0}(X\setminus\{y\})u)$ has stable rank one, and as it is a centrally large subalgebra, this passes to $C(X)\rtimes_{\alpha}\mathbb{Z}$. The RSH constructions in this paper will not be DSH algebra in general. This follows from [3, Corollary 2.22]. Nevertheless, our RSH constructions should be amenable to similar techniques. The question of stable rank of $\mathrm{C}^{*}$-algebras associated to homeomorphisms twisted by (non-trivial) line bundles is the subject of future investigations.